Compute for the value of the acceleration due to gravity g of an object at an altitude equal to twice the radius of the Earth? (radius of Earth = 6.4 x 106 m) 


1
Expert's answer
2021-05-03T07:10:07-0400

"g(0)=G\frac{M}{R^2}"


"g(h)=G\frac{M}{(R+h)^2}" . So, we have


"g(h)=g(0)\cdot (\frac{R}{R+h})^2=g(0)\cdot (\frac{R}{R+R})^2=g(0)/4=9.18/4=2.45\ (m/s^2)" . Answer